# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ C.3 Construction of $A_{n}$
$\left({(C_{\mathbb{B}}\epsilon_{n})^{2}\over 4n^{2p}t_{n}\sigma_{\beta}^{2}}\right)^{t_{n}\over 2}\geq\exp(-d_{6}n\epsilon_{n}^{2})$
@@
C.6 Proof of Condition (C.5)
We will check Condition (C.5) with the choice $C_{1}>(2d_{1}+d_{\sigma}+2)/C_{*}$, where $d_{1}$ is the constant satisfying Condition (C.4). Note that we have
@@
Case 1. Upper bound of $\pi\{T>M_{n}\}$. We will show that $\pi\{T>M_{n}\}e^{(2d_{1}+d_{\sigma}+2)n\epsilon_{n}^{2}}\rightarrow 0$ as $n\rightarrow\infty,$ where $M_{n}=\lfloor C_{1}{n\epsilon_{n}^{2}\over\log n}\rfloor,$ which holds because
@@
$={{1\over n^{(M_{n}+1)C_{*}}}\big{(}1-{1\over n^{(T_{\max}+1)C_{*}}}\big{)}\over 1-{1\over n}}\times{1-{1\over n}\over\big{(}1-{1\over n^{(T_{\max}+1)C_{*}}}\big{)}}$
@@
Hence, $\pi\{T>M_{n}\}e^{(2d_{1}+d_{\sigma}+2)n\epsilon_{n}^{2}}\rightarrow 0$ as $n\rightarrow\infty$. Case 2. Upper bound of $\pi\{\exists t\in[T]\>\text{s.t}\>|\beta_{t}|>n|T\leq M_{n}\}$. We have
@@
where $\beta_{1}\sim N(0,\sigma_{\beta}^{2})$ and $\sigma_{\beta}^{2}>0$ is a constant. Hence, $\pi\{\exists t\in[T]\>\text{s.t}\>|\beta_{t}|>n|T\leq M_{n}\}e^{(2d_{1}+d_{\sigma}+2)n\epsilon_{n}^{2}}\xrightarrow{}0$ as $n\xrightarrow{}\infty$.
C.7 Verification of the conditions in Theorem 4 of nonp1
For given $f\in\mathcal{F}$ and $\sigma^{2}\in\mathbb{R}_{+},$ consider the probability model $Y|\mathbf{X}=f(\mathbf{X})+\epsilon$ with $\epsilon\sim N(0,\sigma^{2})$ and $\mathbf{X}\sim\mathbb{P}_{\mathbf{X}}$. Let $\theta=(f,\sigma^{2})$ and let $p_{\theta}$ be the density of $(\mathbf{X},Y)$ with parameter $\theta$. For given data $\mathbf{x}^{(n)}$ and $\theta=(f,\sigma^{2})$, let $p_{\theta,i}$ be the density of Gaussian distribution $N(f(\mathbf{x}_{i}),\sigma^{2})$ for i=1,,ni=1,...,n.
For given two densities $p_{\theta_{1}},p_{\theta_{2}}$, let $K(p_{\theta_{1}},p_{\theta_{2}})$ be a Kullback-Leibler (KL) divergence defined as $K(p_{\theta_{1}},p_{\theta_{2}})=\int\log(p_{\theta_{1}}(\mathbb{v})/p_{\theta_{2}}(\mathbb{v}))p_{\theta_{1}}(\mathbb{v})d\mathbb{v},$ where $\mathbb{v}\in\mathcal{X}\times\mathbb{R}.$ Let $V(p_{\theta_{1}},p_{\theta_{2}})=\int|\log(p_{\theta_{1}}(\mathbb{v})/p_{\theta_{2}}(\mathbb{v}))-K(p_{\theta_{1}},p_{\theta_{2}})|^{2}p_{\theta_{1}}(\mathbb{v})d\mathbb{v}.$
Let $\mathcal{F}_{\xi}=\{f\in\mathcal{F}:\|f\|_{\infty}\leq\xi\}$ and let $\Theta_{\xi}=\mathcal{F}_{\xi}\times\mathbb{R}_{+}.$ Similarly, we define $\mathcal{F}_{\xi}^{n}=\{f\in\mathcal{F}^{n}:\|f\|_{\infty}\leq\xi\}$ and $\Theta_{\xi}^{n}=\mathcal{F}_{\xi}^{n}\times(1/\xi,\xi).$ Following Theorem 4 of nonp1 , to show the posterior concentration rate in (C.2), it suffices to verify that for all $f_{0}\in\mathcal{H}^{\alpha}_{0,F}$ and $\mathbf{x}^{(n)}\in A_{n}$, there exists a sieve $\Theta_{\xi}^{n}$ and a constant $d_{*}>0$ such that the following three conditions are satisfied.
$\log N\left({\epsilon_{n}\over 36},\Theta^{n}_{\xi},\|\cdot\|_{2,n}\right)\leq n\epsilon_{n}^{2}$ (C.16)
$\pi_{\xi}\{\mathbb{B}_{n}\}\geq e^{-d_{*}n\epsilon_{n}^{2}}$ (C.17)
$\pi_{\xi}\{\Theta_{\xi}\backslash\Theta^{n}_{\xi}\}\leq e^{-(d_{*}+2)n\epsilon_{n}^{2}},$ (C.18)
@@ C.7 Verification of the conditions in Theorem 4 of nonp1
$\mathbb{B}_{n}:=\bigg{\{}\theta\in\Theta_{\xi}:{1\over n}\sum_{i=1}^{n}K(p_{\theta_{0},i},p_{\theta,i})\leq\epsilon_{n}^{2},{1\over n}\sum_{i=1}^{n}V(p_{\theta_{0,i}},p_{\theta,i})\leq\epsilon_{n}^{2}\bigg{\}}.$
We will show that the above three conditions are satisfied to complete the proof of Theorem 5.1. Verification of Condition (C.16). Note that
$N\left({\epsilon_{n}\over 36},\Theta^{n}_{\xi},\|\cdot\|_{2,n}\right)\leq N\left(\epsilon_{n}/72,\mathcal{F}^{n}_{\xi},\|\cdot\|_{\infty}\right)N\left(\epsilon_{n}/72,(1/\xi,\xi),|\cdot|\right).$
Since $N\left(\epsilon_{n}/72,(1/\xi,\xi),|\cdot|\right)\leq\left(1+{72\max(1/\xi,\xi)\over\epsilon_{n}}\right)$, Condition (C.3) implies
$\log N\left({\epsilon_{n}\over 36},\Theta^{n}_{\xi},\|\cdot\|_{2,n}\right)\lesssim n\epsilon_{n}^{2}.$
Verification of Condition (C.17). Direct calculation yields,
${1\over n}\sum_{i=1}^{n}K(p_{\theta_{0},i},p_{\theta,i})$ $={1\over 2}\log\bigg{(}{\sigma^{2}\over\sigma_{0}^{2}}\bigg{)}-{1\over 2}\bigg{(}1-{\sigma_{0}^{2}\over\sigma^{2}}\bigg{)}+{\|f-f_{0}\|_{2,n}^{2}\over 2\sigma^{2}},$
${1\over n}\sum_{i=1}^{n}V(p_{\theta_{0},i},p_{\theta,i})$ $={1\over 2}\bigg{(}1-{\sigma_{0}^{2}\over\sigma^{2}}\bigg{)}^{2}+{\sigma_{0}^{2}\|f-f_{0}\|_{2,n}^{2}\over\sigma^{2}}.$
As in the proof of Theorem 2 in artbart , by using Taylor expansion, there exist a positive constant $C_{\mathbb{B}}$ such that
$\mathbb{B}_{n}\supseteq\{\theta\in\Theta_{\xi}:\|f-f_{0}\|_{2,n}\leq C_{\mathbb{B}}\epsilon_{n},|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}}\epsilon_{n}\}.$
Additionally, we have
$\pi_{\xi}\{\theta\in\Theta_{\xi}:\|f-f_{0}\|_{2,n}\leq C_{\mathbb{B}}\epsilon_{n},|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}}\epsilon_{n}\}$
$\geq\pi_{\xi}\{\theta\in\Theta_{\xi}:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n},|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}}\epsilon_{n}\}.$
In turn, since $\xi\geq\max\{2^{p}F,1/\sigma^{2}_{\min},\sigma^{2}_{\max}\}+C_{\mathbb{B}}\epsilon_{n}$ for sufficiently large $n$, we have
$\pi_{\xi}\{\theta\in\Theta_{\xi}:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n},|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}}\epsilon_{n}\}$
$\geq\pi\{f:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}\}\pi\{\sigma^{2}:|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}}\epsilon_{n}\},$
and from the proof of Theorem 2 in artbart , we have $\pi\{\sigma^{2}:|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}}\epsilon_{n}\}\gtrsim{1\over n}$. Let $d_{\sigma}$ be a positive constant such that $\pi\{\sigma^{2}:|\sigma^{2}-\sigma_{0}^{2}|\leq C_{\mathbb{B}}\epsilon_{n}\}\geq\exp(-d_{\sigma}n\epsilon_{n}^{2})$. Thus, Condition (C.4) implies
$\pi_{\xi}\bigg{\{}\theta\in\Theta_{\xi}:{1\over n}\sum_{i=1}^{n}K(p_{\theta_{0},i},p_{\theta,i})\leq\epsilon_{n}^{2},{1\over n}\sum_{i=1}^{n}V(p_{\theta_{0,i}},p_{\theta,i})\leq\epsilon_{n}^{2}\bigg{\}}\geq e^{-(d_{1}+d_{\sigma})n\epsilon_{n}^{2}}.$
@@
which is less than $e^{-(d_{1}+d_{\sigma}+2)n\epsilon_{n}^{2}}$ by Condition (C.4) and Condition (C.5) whenever $2^{p}F+C_{\mathbb{B}}\epsilon_{n}\leq\xi$. The proof is completed by letting $d_{*}=d_{1}+d_{\sigma}$. ∎
Appendix D Bridging the empirical and populational identifiabilities for multinary-product trees
For any component $S\subseteq[p]$, we first define a multinary-product tree, which is an extension of a binary-product tree, as follows. Recall that $\mathcal{X}_{j}=[0,1]$ for all $j\in[p].$ A partition $\mathcal{P}_{j}$ of $\mathcal{X}_{j}$ is called an interval partition if any element in $\mathcal{P}_{j}$ is an interval. For example, $\{[0,1/3),[1/3,2/3),[2/3,1]\}$ is an interval partition. In the followings, we only consider interval partitions and we simply write them as ‘partitions’ unless there is any confusion. Let $\mathcal{P}_{j}$ be a given partition of $\mathcal{X}_{j}$ for each $j\in S,$ where $\phi_{j}:=|\mathcal{P}_{j}|\geq 2.$ Let $\mathcal{P}=\prod_{j\in S}\mathcal{P}_{j}.$ Then, the multinary-product tree defined on the product partition $\mathcal{P}$ with the height vector $\boldsymbol{\gamma}=(\gamma_{\boldsymbol{\ell}},\boldsymbol{\ell}\in\prod_{j\in S}\{1,\ldots,\phi_{j}\})$ is defined as
@@ Appendix D Bridging the empirical and populational identifiabilities for multinary-product trees
where $I_{\boldsymbol{\ell}}=\prod_{j\in S}I_{j,\ell_{j}},I_{j,\ell_{j}}\in\mathcal{P}_{j}$. Note that a binary-product tree is a special case of a multinary-product tree with $\phi_{j}=2$ for all $j\in S.$ Figure D.3 compares binary-product tree and multinary-product tree.
Figure D.3: Examples of binary-product tree and multinary-product tree. The left panel is a binary-product tree, while the right panel illustrates the partition and cell heights of a multinary-product tree with $\phi_{1}=5$ and $\phi_{2}=3$.
D.1 Identifiable transformation of multinary-product tree

### v3

@@ B.3 Construction of $A_{n}$
$\left({(C_{\mathbb{B}}\epsilon_{n})^{2}\over 4(1+n)^{2p}t_{n}\sigma_{\beta}^{2}}\right)^{t_{n}\over 2}\geq\exp(-d_{6}n\epsilon_{n}^{2})$
@@
B.6 Proof of Condition (B.5)
We will check Condition (B.5) with the choice $C_{1}>(2d_{1}+2)/C_{*}$, where $d_{1}$ is the constant satisfying Condition (B.4). Note that we have
@@
Case 1. Upper bound of $\pi\{T>M_{n}\}$. We will show that $\pi\{T>M_{n}\}e^{(2d_{1}+2)n\epsilon_{n}^{2}}\rightarrow 0$ as $n\rightarrow\infty,$ where $M_{n}=\lfloor C_{1}{n\epsilon_{n}^{2}\over\log n}\rfloor,$ which holds because
@@
$\leq{{1\over n^{(M_{n}+1)C_{*}}}\big(1-{1\over n^{(T_{\max}+1)C_{*}}}\big)\over 1-{1\over n^{C_{*}}}}\times{1-{1\over n^{C_{*}}}\over\big(1-{1\over n^{(T_{\max}+1)C_{*}}}\big)}$
@@
Hence, $\pi\{T>M_{n}\}e^{(2d_{1}+2)n\epsilon_{n}^{2}}\rightarrow 0$ as $n\rightarrow\infty$. Case 2. Upper bound of $\pi\{\exists t\in[T]\>\text{s.t}\>|\beta_{t}|>n|T\leq M_{n}\}$. We have
@@
where $\beta_{1}\sim N(0,\sigma_{\beta}^{2})$ and $\sigma_{\beta}^{2}>0$ is a constant. Hence, $\pi\{\exists t\in[T]\>\text{s.t}\>|\beta_{t}|>n|T\leq M_{n}\}e^{(2d_{1}+2)n\epsilon_{n}^{2}}\xrightarrow{}0$ as $n\xrightarrow{}\infty$.
B.7 Verification of the conditions in Theorem 4 of Ghosal and van der Vaart (2007)
For a given $f\in\mathcal{F}$, consider the probability model $Y|\mathbf{X}\sim\mathbb{P}_{f(\mathbf{X}),1}$ and $\mathbf{X}\sim\mathbb{P}_{\mathbf{X}}$. For given data $\mathbf{x}^{(n)}$ and $f$, let $p_{f(\mathbf{x}_{i})}$ be the density function of $\mathbb{P}_{f(\mathbf{x}_{i}),1}$ with respect to Lebesgue measure for i=1,,ni=1,...,n. For given two densities $p_{f_{1}},p_{f_{2}}$, let $K(p_{f_{1}},p_{f_{2}})$ be a Kullback-Leibler (KL) divergence defined as $K(p_{f_{1}},p_{f_{2}})=\int\log(p_{f_{1}}(\mathbb{v})/p_{f_{2}}(\mathbb{v}))p_{f_{1}}(\mathbb{v})d\mathbb{v},$ where $\mathbb{v}\in\mathcal{X}\times\mathbb{R}.$ Let $V(p_{f_{1}},p_{f_{2}})=\int|\log(p_{f_{1}}(\mathbb{v})/p_{f_{2}}(\mathbb{v}))-K(p_{f_{1}},p_{f_{2}})|^{2}p_{f_{1}}(\mathbb{v})d\mathbb{v}.$ For i=1,,ni=1,...,n, let $p_{f_{1}(\mathbf{x}_{i})}$, $p_{f_{2}(\mathbf{x}_{i})}$ be densities and let $\rho_{n}^{2}(\mathbb{P}_{f_{1},1},\mathbb{P}_{f_{2},1})={1\over n}\sum_{i=1}^{n}\int(\sqrt{p_{f_{1}(\mathbf{x}_{i})}(y_{i})}-\sqrt{p_{f_{2}(\mathbf{x}_{i})}(y_{i})})^{2}dy_{i}$. Let $\mathcal{F}_{\xi}=\{f\in\mathcal{F}:\|f\|_{\infty}\leq\xi\}$ and $\mathcal{F}_{\xi}^{n}=\{f\in\mathcal{F}^{n}:\|f\|_{\infty}\leq\xi\}$. Note that by Theorem 7.6 in page 81 of Ibragimov and Has’ Minskii (1983), for all sufficiently large $n$ and any $f\in\mathcal{F}_{\xi}$,
$C_{\rho,1}\|f_{0}-f\|_{2,n}\leq\rho_{n}(\mathbb{P}_{f_{0},1},\mathbb{P}_{f,1})\leq C_{\rho,2}\|f_{0}-f\|_{2,n}.$ (B.17)
holds, where $C_{\rho,1}$ and $C_{\rho,2}$ are positive constants. Therefore, to show the posterior concentration rate in (B.2), following Theorem 4 of Ghosal and van der Vaart (2007) and (B.17), it suffices to verify that for all $f_{0}\in\mathcal{H}^{\alpha}_{0,F}$ and $\mathbf{x}^{(n)}\in A_{n}$, there exists a sieve $\mathcal{F}_{\xi}^{n}$ and a constant $d_{*}>0$ such that the following three conditions are satisfied.
$\log N\left({\epsilon_{n}\over 36},\mathcal{F}^{n}_{\xi},\|\cdot\|_{2,n}\right)\leq n\epsilon_{n}^{2}$ (B.18)
$\pi_{\xi}\{\mathbb{B}_{n}\}\geq e^{-d_{*}n\epsilon_{n}^{2}}$ (B.19)
$\pi_{\xi}\{\mathcal{F}_{\xi}\backslash\mathcal{F}^{n}_{\xi}\}\leq e^{-(d_{*}+2)n\epsilon_{n}^{2}},$ (B.20)
@@ B.7 Verification of the conditions in Theorem 4 of Ghosal and van der Vaart (2007)
$\mathbb{B}_{n}:=\bigg\{f\in\mathcal{F}_{\xi}:{1\over n}\sum_{i=1}^{n}K(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})\leq\epsilon_{n}^{2},{1\over n}\sum_{i=1}^{n}V(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})\leq\epsilon_{n}^{2}\bigg\}.$
We will show that the above three conditions are satisfied to complete the proof of Theorem 5.1. Verification of Condition (B.18). Condition (B.3) implies
$\log N\left({\epsilon_{n}\over 36},\mathcal{F}^{n}_{\xi},\|\cdot\|_{2,n}\right)\lesssim n\epsilon_{n}^{2}.$
Verification of Condition (B.19). Direct calculation yields,
$K(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})$ $=\int\Big((f_{0}(\mathbf{x}_{i})-f(\mathbf{x}_{i}))y-A(f_{0}(\mathbf{x}_{i}))+A(f(\mathbf{x}_{i}))\Big)p_{f_{0}(\mathbf{x}_{i})}(y)dy$ (B.21)
$=\bigg((f_{0}(\mathbf{x}_{i})-f(\mathbf{x}_{i}))\mathbb{E}[Y_{i}|\mathbf{x}_{i}]-A(f_{0}(\mathbf{x}_{i}))+A(f(\mathbf{x}_{i}))\bigg)$ (B.22)
$=\bigg((f_{0}(\mathbf{x}_{i})-f(\mathbf{x}_{i}))\dot{A}(f_{0}(\mathbf{x}_{i}))-A(f_{0}(\mathbf{x}_{i}))+A(f(\mathbf{x}_{i}))\bigg)$ (B.23)
for i=1,,ni=1,...,n, and
$V(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})$ $=\int(f_{0}(\mathbf{x}_{i})-f(\mathbf{x}_{i}))^{2}(y-\dot{A}(f_{0}(\mathbf{x}_{i})))^{2}p_{f_{0}(\mathbf{x}_{i})}(y)dy$ (B.24)
$=(f_{0}(\mathbf{x}_{i})-f(\mathbf{x}_{i}))^{2}Var(Y_{i}|\mathbf{x}_{i})$ (B.25)
$=(f_{0}(\mathbf{x}_{i})-f(\mathbf{x}_{i}))^{2}\ddot{A}(f_{0}(\mathbf{x}_{i})).$ (B.26)
In turn, using Talyor expansion, we have
$K(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})={1\over 2}\ddot{A}(\tilde{x})(f_{0}(\mathbf{x}_{i})-f(\mathbf{x}_{i}))^{2},$
where $\tilde{x}\in[-F,F]$.
That is, we have
$\max\bigg\{{1\over n}\sum_{i=1}^{n}K(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})}),{1\over n}\sum_{i=1}^{n}V(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})\bigg\}\leq C_{A}\|f_{0}-f\|_{2,n}^{2}.$
With a positive constant $C_{\mathbb{B}}=\sqrt{C_{A}}$, we have
$\mathbb{B}_{n}\supseteq\{f\in\mathcal{F}_{\xi}:\|f-f_{0}\|_{2,n}\leq C_{\mathbb{B}}\epsilon_{n}\}$
and
$\pi_{\xi}\{f\in\mathcal{F}_{\xi}:\|f-f_{0}\|_{2,n}\leq C_{\mathbb{B}}\epsilon_{n}\}\geq\pi_{\xi}\{f\in\mathcal{F}_{\xi}:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}\}.$
In turn, since $\xi\geq 2^{p}F+C_{\mathbb{B}}\epsilon_{n}$ for sufficiently large $n$, we have
$\pi_{\xi}\{f\in\mathcal{F}_{\xi}:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}\}\geq\pi\{f\in\mathcal{F}:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}\}.$
Thus, Condition (B.4) implies
$\pi_{\xi}\bigg\{f\in\mathcal{F}_{\xi}:{1\over n}\sum_{i=1}^{n}K(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})\leq\epsilon_{n}^{2},{1\over n}\sum_{i=1}^{n}V(p_{f_{0}(\mathbf{x}_{i})},p_{f(\mathbf{x}_{i})})\leq\epsilon_{n}^{2}\bigg\}\geq e^{-d_{1}n\epsilon_{n}^{2}}.$
@@
which is less than $e^{-(d_{1}+2)n\epsilon_{n}^{2}}$ by Condition (B.4) and Condition (B.5) whenever $2^{p}F+C_{\mathbb{B}}\epsilon_{n}\leq\xi$. The proof is completed by letting $d_{*}=d_{1}$. ∎
Appendix C Bridging the empirical and populational identifiabilities for multinary-product trees
For any component $S\in\mathbb{S}$, we first define a multinary-product tree, which is an extension of a binary-product tree, as follows. Recall that $\mathcal{X}_{j}=[0,1]$ for all $j\in[p].$ A partition $\mathcal{P}_{j}$ of $\mathcal{X}_{j}$ is called an interval partition if any element in $\mathcal{P}_{j}$ is an interval. For example, $\{[0,1/3),[1/3,2/3),[2/3,1]\}$ is an interval partition. In the followings, we only consider interval partitions and we simply write them as ‘partitions’ unless there is any confusion. Let $\mathcal{P}_{j}$ be a given partition of $\mathcal{X}_{j}$ for each $j\in S,$ where $\phi_{j}:=|\mathcal{P}_{j}|\geq 2.$ Let $\mathcal{P}=\prod_{j\in S}\mathcal{P}_{j}.$ Then, the multinary-product tree defined on the product partition $\mathcal{P}$ with the height vector $\boldsymbol{\gamma}=(\gamma_{\boldsymbol{\ell}},\boldsymbol{\ell}\in\prod_{j\in S}\{1,\ldots,\phi_{j}\})$ is defined as
@@ Appendix C Bridging the empirical and populational identifiabilities for multinary-product trees
where $I_{\boldsymbol{\ell}}=\prod_{j\in S}I_{j,\ell_{j}},I_{j,\ell_{j}}\in\mathcal{P}_{j}$. Note that a binary-product tree is a special case of a multinary-product tree with $\phi_{j}=2$ for all $j\in S.$ Figure C.4 compares binary-product tree and multinary-product tree.
Figure C.4: Examples of binary-product tree and multinary-product tree. The left panel is a binary-product tree, while the right panel illustrates the partition and cell heights of a multinary-product tree with $\phi_{1}=5$ and $\phi_{2}=3$.
C.1 Identifiable transformation of multinary-product tree